Tracy is borrowing $10,800 on a six-year, 15%, add-on interest loan. What will Tracy's monthly payments be?
Answer #1

Interest on Loan = 10,800(0.15)(6) = $9,720

Total Payment = 10,800 + 9,720 = $20,520

Monthly Payment = 20,520/72 = $285